OpenStudy (anonymous):

A trapezoid is drawn on a grid with vertices at P(-5, 7), Q(-2, 7), R(2, 3), and S(-8, 3) . Trapezoid UVWX will be drawn similar to Trapezoid PQRS with side PQ corresponding to side UV. The point (-2, -1) will be the location of vertex U and point (7, -1) will be the location of point V. What will be the coordinates of vertex X? Show all your work or explain your reasoning.

OpenStudy (anonymous):

this is as far as i got To find the vertex for X you must use the proportion UV/PQ=UX/PS. That amounts to 9/3 = 15/5. 9/3 and 15/5 equal 3 so the scale factor is 3:1.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

you there isuri

OpenStudy (anonymous):

*suru

OpenStudy (anonymous):

***[ISURU]*** can you help me

OpenStudy (anonymous):

im here

OpenStudy (yttrium):

You said, you have scale factor of 3:1. Therefore, you can then find the dimensions of the other sides of the trapezoid UVWX with traperzoid PQRS as reference.

OpenStudy (anonymous):

-24,9

OpenStudy (yttrium):

May I see your solution?

OpenStudy (anonymous):

-8 * 3=-24 and 3*3=9

OpenStudy (yttrium):

what is -24,9 ?

OpenStudy (anonymous):

X

OpenStudy (anonymous):

is my answer right??????

OpenStudy (yttrium):

I'm not sure with your solution because this wouldn't be that easy. How did you get the distance UV and PQ? May I know?

OpenStudy (anonymous):

distance formula

OpenStudy (yttrium):

did you already find the distance PQ and RS?

OpenStudy (anonymous):

no

OpenStudy (yttrium):

I think you need to find first the distance of WX using the scale ratio of 3 which must be based from the distance of RS and you can find RS using distance formula, right?

OpenStudy (anonymous):

i guess

OpenStudy (yttrium):

Are you done doing it now?

OpenStudy (anonymous):

RS is 10

OpenStudy (yttrium):

Then WX is?

OpenStudy (anonymous):

i dont know because X is unknown

OpenStudy (yttrium):

isn't it that you can use the relation of similar figures right? because it is written in the problem that those trapezoids are similar.

OpenStudy (anonymous):

right

OpenStudy (yttrium):

So WX is now then?

OpenStudy (anonymous):

i really dont know

OpenStudy (yttrium):

You said the the scale ratio of UVWX to PQRS is 3:1. Therefore, if RS = 10, WX = 30. You get the point?

OpenStudy (anonymous):

so the lengths are just mmultiplied by 3

OpenStudy (yttrium):

Yes. :) Can you find too the distance of XU?

OpenStudy (yttrium):

@JackJones

OpenStudy (anonymous):

how would you do that

OpenStudy (yttrium):

Find PS and multiply it by 3. :)

OpenStudy (anonymous):

PS is 5 so 15

OpenStudy (yttrium):

Yes. Now you can use system of equality to solve for the coordinates of X using the computed distance and the distance formula for WX and UX.

OpenStudy (anonymous):

could you do that part for me plz

OpenStudy (anonymous):

ganshie plz help